# XiYan-SQL: A Novel Multi-Generator Framework For Text-to-SQL

### v1

@@ Abstract
Index Terms:
Generation, Text-to-SQL, Multiple model ensemble, LLM fine-tuning strategy.
11footnotetext: Preprint. Under review.
@@ I Introduction
The task of translating natural language (NL) queries into structured query language (SQL) queries, known as Text-to-SQL, or NL2SQL, is a long-standing task for the development of NL interfaces to relational database management systems. This capability significantly reduces the interaction cost of users accessing database systems. Recent research is empowered by advances in large language models (LLMs), which have significantly improved the performance of NL2SQL solutions [1]. A series of approaches based on in-context learning (ICL) have been developed to enhance the SQL generation performance of these models by designing different prompts, such as chain-of-thought (CoT) [2], question decomposition [3], and choice of demonstration examples [4]. Nonetheless, these methods are highly constrained by the sensitivity of LLMs to the structure and content of prompts.
@@
Consequently, an established effective strategy is to generate multiple SQL candidates by designing various forms of prompts, followed by an SQL selection process. Compared to a single SQL query, this approach [4, 5] of utilizing multiple forms of prompts enhances the diversity of the candidate samples and demonstrates promising performance on the challenging BIRD benchmark [6]. However, relying solely on the design of multiple prompts does not adequately ensure the quality and robustness of SQL generation while improving candidate diversity. And utilizing self-consistency as a selection criterion does not always yield the optimal SQL query.
@@
To address these challenges, we propose XiYan-SQL, an end-to-end SQL generation framework, which focuses on generating SQL candidates with high accuracy and diversity, as well as effectively selecting the best candidate. We first introduce a Schema Filter Module that provides SQL generators with diverse and high-quality database schemas. This module performs multi-path retrieval of tables, columns, and values based on the inputs and subsequently employs a model for iterative column selection.
@@
For the generation of SQL candidates, we propose a novel method that integrates multiple generators, leveraging the distinct advantages of each to generate high-quality and diverse SQL queries. Instead of designing different forms of prompts, we capitalize on the high controllability of supervised fine-tuning (SFT) to construct multiple SQL generation models. Specifically, inspired by the multi-tasking approach in natural language processing, we explore a multi-task joint fine-tuning strategy tailored for SQL generation. This strategy involves the incorporation of various SQL and natural language transformation tasks to enhance the model’s syntactic alignment capabilities. Additionally, we investigate the multi-format SQL enhancement training approach aimed at developing generators with distinct generation advantages to improve the diversity of candidates. Ultimately, these strategies significantly improve the performance of fine-tuned models.
@@
For the SQL candidate selection phase, we employ a selection model, fine-tuned using constructed comparative samples, to identify the best answer among all candidates. Furthermore, we explore a candidate reorganization strategy based on the consistency results, to improve the selection model’s attention to potentially correct results, thereby improving selection performance.
@@
In the experiments, we evaluate XiYan-SQL on two benchmarks, BIRD [6] and Spider [7], to demonstrate the superiority of the framework. For the well-known and challenging BIRD benchmark, we achieve an execution accuracy of 75.63%, surpassing all other methods and establishing a new SOTA performance. For the Spider test set, we achieve SOTA performance with an execution accuracy of 89.65%. The extensive experimental results and comprehensive analysis further confirm the effectiveness and robustness of our method. The related code and models have been gradually released to support community research. ***https://github.com/XGenerationLab/XiYan-SQL
Our main contributions are as follows:
@@
  • We propose a candidate reorganization strategy that utilizes an SQL selection model to identify the optimal SQL from multiple candidates.
@@ II Related Work
Prompt Engineering Methods.    Prompt engineering methods leverage the inherent capabilities of models to guide LLMs in generating diverse SQL queries by optimizing prompts [18, 13, 5, 14]. These methods are typically employed out-of-the-box on closed-source models with a vast number of parameters, such as GPT-4[19] and Gemini[20]. Some approaches based on in-context learning address various stages of the workflow, including schema linking[21, 22], SQL generation [1], and SQL refinement[14, 3, 23], by designing question/task decomposition[3], demonstration selection[4], chain-of-thought reasoning[2], and multi-agent processing within the prompts. To enhance performance, an effective strategy is to generate multiple candidate SQL queries through different prompt designs and subsequently select the best one [4, 5]. However, these methods rely solely on prompt design, which presents challenges related to the sensitivity of LLMs to prompts, potentially compromising the quality and robustness of SQL generation, along with incurring significant inference overhead. In contrast, our approach effectively mitigates these problems by constructing different high-accuracy fine-tuned SQL generation models.
@@
This paper focuses on the development of high-accuracy SQL generation models. By utilizing the proposed multi-task fine-tuning strategy and multi-format data manner, the fine-tuned models demonstrate superior performance compared to the ICL-based closed-source model. Furthermore, to the best of our knowledge, this work represents the first exploration in this field to generate multiple candidate SQL queries through multi-generators. Through the comprehensive XiYan-SQL framework, we can effectively generate multiple candidates and select the optimal one.
@@ III-A Overall Framework
We present the XiYan-SQL framework in Figure 1, where the inputs include a question, evidence (i.e., external knowledge), and the database schema (i.e., the overall organization of tables, columns, and values within the database). The framework consists of three steps: Schema Filter, Multiple SQL Generation, and SQL Selection. The Schema Filter module initially generates multiple filtered schemas. Subsequently, multiple candidate SQL queries are generated through various SQL generators. Finally, the most accurate SQL query is selected by the selection model based on the candidate reorganization. The details of each step are described below.
@@ III-B Schema Filter
Multi-path Retrieval.   Multi-path Retrieval is a pruning mechanism for large databases that employ cosine similarity for the embeddings of all tables, columns, and values. Initially, we used an out-of-the-box LLM $\mathcal{M}_{s}$ to extract keywords from the question $Q$ and evidence $E$, represented as $K=f_{\mathcal{M}_{s}}(Q,E)$. These keywords, denoted as $K=\{k_{1},k_{2},\ldots\}$, are instrumental in identifying relevant columns and values within the database. In the table and column retrieval stage, we compute the scores between the keywords $K$ and the original schema $S$, which can be represented as a set of columns, $\{c_{1},c_{2},\ldots\}$. This process is calculated as the product of two distinct scores as follows.
@@
Column Selection.   Column Selection introduces diversity into the retrieved schemas, facilitating subsequent multiple SQL generation. This process is detailed in Algorithm 1, which outputs $p_{s}$ different schemas through multiple iterations. In each iteration, we prompt an LLM $\mathcal{M}_{s}$ to select columns related to the question and evidence, saved as $S^{slct}_{i}$. The function PFKeyIdentifier then identifies primary and foreign keys according to the columns in $S^{slct}_{i}$. A new schema is generated by unifying the previous schemas with the selected columns. The selected columns in this iteration are removed from $S^{rtrv}$, while retaining essential columns such as primary keys. As iterations increase, the precision of the selected schemas tends to decrease while recall improves, resulting in a diversity of filtered schemas.
@@
7:     $S^{rtrv}\leftarrow S^{rtrv}\setminus(S^{slct}_{i}\setminus P_{i})$Remove the selected}
@@ III-C Multiple SQL Generation
Due to the inherent flexibility of natural language forms, this unidirectional conversion method is insufficient to fully exploit the model’s semantic linking capabilities. To address this issue and enhance the semantic alignment capability of LLMs for this task, we further design auxiliary alignment tasks and explicitly fine-tune the SQL generation model for execution. In response to natural language questions, we design a reverse inference task, as illustrated in Figure 2(b), which aims to encourage the model to infer a set of potential questions based on SQL and relevant information. For natural language evidence knowledge, we establish a task focused on the reverse inference of evidence, shown in Figure 2(c). This task involves the model identifying the most relevant evidence from a set of candidate evidence based on the SQL generation process. Additionally, for SQL queries, we develop a self-refine task for the model, as depicted in Figure 2(d), which entails regenerating SQL after refinement based on previous SQL generation processes and execution results. By leveraging a series of specialized tasks to jointly fine-tune models, our developed single generator surpasses ICL-based closed-source models, such as GPT-4o and Gemini, resulting in the generation of higher-quality SQL queries.
Figure 2: The illustration of the process for multiple related tasks.
We further develop from a single generator to multiple generators in order to enhance the diversity of generated candidates. For a single user question, there can be multiple SQL queries that correspond to fulfill the intent. Therefore, we aim to develop additional models with different ways of generation to increase the likelihood of arriving at the correct answers through diverse outputs. To achieve this, we restructure the format variations with SQL to introduce greater diversity during model building. A key aspect is the variation of SQL writing patterns, encouraging the model to attempt using more complex SQL structures, thereby enhancing its ability to cover a wider range of challenging scenarios (as illustrated by $\mathrm{SQL}_{1}$ on the left side of Figure 3, which demonstrates a preference for utilizing advanced queries). Furthermore, we also implement certain modifications to the aspect of SQL writing style, allowing the model to adopt variations in its writing conventions to increase diversity in the generated outputs (as demonstrated on the right side of Figure 3, $\mathrm{SQL}_{3}$ presents a more standardized output representation). In practice, we conduct various aspects of reformulation on a subset of challenging queries to supplement the diverse training data. By leveraging these formatted and multi-task data for joint fine-tuning, we obtain high-accuracy models with diverse SQL generation preferences.
@@
Figure 3: Examples of multiple SQL queries with different format corresponding to the same input. The middle SQL query is a typical response. In contrast, the left SQL employs a chunked writing pattern, which exhibits greater complexity in structure compared to the middle; the SQL on the right differs from that in the middle in terms of presentation style.
@@ III-D SQL Selection
After obtaining multiple SQL candidates, the challenge remains to select the most accurate and reasonable SQL query from these candidates. Most candidate selection methods [4, 21] employ self-consistency [33] to select the SQL query that appears most consistently across multiple candidate samples. However, there are limitations present: it cannot adequately handle situations where there is a lack of majority consensus, and even the most consistent result is not always the correct case. To this end, we propose an enhanced selection strategy that employs a selection model to make decisions based on consistent results.
To develop a selection model, we focus on constructing enhanced synthetic training data. On the one hand, we employ various generators with diverse inputs to generate multiple candidate SQL queries. By grouping and clustering samples based on execution results, we are able to construct positive and negative contrastive samples. Additionally, to account for the complexity of the candidate results, we utilize an LLM to make a few controlled modifications for the correct queries to supplement the dataset. On the other hand, when organizing the candidate results into training prompts, we first perform a de-formalization process on the samples to reduce the interference caused by the stylistic aspects of SQL. For the constructed dataset, we ensure that the distribution of the combinations of positive and negative samples is balanced, that the generators from which the selected candidates originate are evenly distributed, and that the order of the candidates is uniformly distributed.
Due to the presence of numerous candidate information in the model selection process, we further propose a candidate reorganization strategy to enhance the model’s attention. The complete selection process is illustrated in Algorithm 3. Initially, we obtain the execution results $R=\{r_{1},\ldots,r_{n}\}$ of the candidate SQL queries $L=\{l_{1},\ldots,l_{n}\}$ on the database, ensuring that any candidates with execution errors have already been excluded. We group the candidates based on the consistency of their execution results to obtain a collection of clusters $\mathcal{C}=\{C_{1},...,C_{m}\}$. Subsequently, we perform inter-group and intra-group sorting. The inter-group sorting is performed in descending order based on the size of each group, resulting in $\mathcal{C}^{\prime}$. The intra-group sorting is executed according to the order of the generators $\mathcal{O}$ corresponding to the candidates, resulting in $\mathcal{C}^{\prime\prime}$, where the sequence $\mathcal{O}$ is arranged in descending order based on the performance evaluation of the generators.
@@
1:  Clustering $\mathcal{C}=\{C_{1},...,C_{m}\}\leftarrow\text{groupby}(L,R)$
2:  Schema $\mathcal{S}^{un}\leftarrow\text{schema\_union}(\mathcal{S})$
3:  if $|\mathcal{C}|=1$ then
4:     return  $l^{*}\leftarrow\arg_{l\in L}\min|l|$All are consistent; the shortest is selected.}
5:  else
6:     Clusters $\mathcal{C}^{\prime}\leftarrow\text{sort}(\mathcal{C})$ by size, descending {Inter-group sorting}
7:     Initialize Clusters $\mathcal{C}^{\prime\prime}\leftarrow\emptyset$
8:     for each cluster $C^{\prime}_{i}\in\mathcal{C}^{\prime}$ do
9:        Append $C^{\prime\prime}_{i}\leftarrow\text{sort}(C^{\prime}_{i})$ order by $\mathcal{O}$Intra-group sorting}
10:     Initialize Candidate $L^{\prime}\leftarrow\emptyset$, indicator $j\leftarrow 0$
11:     if $|C_{1}^{\prime\prime}|\geq\lceil{|L|}/2\rceil$ then
12:        for each Cluster $C_{i}^{\prime\prime}\in\mathcal{C}^{\prime\prime}$ do
13:           for each SQL $l\in C_{i}^{\prime\prime}$ do
14:              Append $L_{j}^{\prime}\leftarrow l$ , $j\leftarrow j+1$
15:     else
16:        for each Cluster $C_{i}^{\prime\prime}\in\mathcal{C}^{\prime\prime}$ do
17:           Append $L_{j}^{\prime}\leftarrow\arg_{l\in C_{i}^{\prime\prime}}\min|l|$, $j\leftarrow j+1$
18:     return  $l^{*}\leftarrow f_{\mathcal{M}_{c}}(Q,\mathcal{S}^{un},E,L^{\prime})$Selection model prediction}
If there exists a consensus result that predominates within the clusters, namely $|C_{1}^{\prime\prime}|\geq\lceil{|L|}/2\rceil$, we organize all candidates sequentially according to order $\mathcal{C}^{\prime\prime}$ and present them to the selection model $\mathcal{M}_{c}$. Conversely, we select one candidate from each group within $\mathcal{C}^{\prime\prime}$ and assemble the results in accordance with the inter-group ordering to present to $\mathcal{M}_{c}$. The model $\mathcal{M}_{c}$ takes into account the question, the union of all filtered schemas, the evidence, and the reorganized candidates to make its decision. Due to the selection tendency of LLMs towards options that appear earlier [34], employing this clustering and ranking approach to organize candidates allows us to effectively leverage strong consistency priors while mitigating the influence of interfering information. This allows the model’s attention to be concentrated, thereby enhancing their comprehension ability.

### v2

@@ I Introduction
The task of translating natural language (NL) queries into structured query language (SQL) queries, known as Text-to-SQL, or NL2SQL, is a long-standing task for the development of NL interfaces to relational database management systems. This capability significantly reduces the interaction cost of users accessing database systems. Recent research is empowered by advances in large language models (LLMs), which have significantly improved the performance of NL2SQL solutions [9]. A series of approaches based on in-context learning (ICL) has been developed to enhance the SQL generation performance of these models by designing different prompts, such as chain-of-thought (CoT) [36], question decomposition [33], and choice of demonstration examples [8]. Nonetheless, these methods are highly constrained by the sensitivity of LLMs to the structure and content of prompts.
@@
Consequently, a commonly adopted strategy is to generate multiple SQL candidates by designing various forms of prompts, followed by an SQL selection process. Compared to a single SQL query, this approach [8, 33] of utilizing multiple forms of prompts enhances the diversity of the candidate samples and improves performance on challenging benchmarks such as BIRD [13]. However, we observe that increased candidate diversity alone does not guarantee correctness or robustness: generated candidates can still contain errors and inconsistencies. Compact, task-specialized fine-tuned SQL models can often provide higher-quality candidates in this setting [11, 24]. Moreover, utilizing self-consistency as a selection criterion does not always yield the optimal SQL query[22].
@@
To address these challenges, we propose XiYan-SQL111“XiYan” is the Chinese name of our ChatBI product., an end-to-end framework built on a novel architectural philosophy: a hybrid ensemble of expert generators. Instead of relying on a single, monolithic model, our approach is centered on a core of specialized, fine-tuned models, augmented by the ICL-based model. This sophisticated integration of multiple, diverse generators allows us to achieve superior accuracy and diversity. We first introduce a Schema Filter Module that provides SQL generators with diverse and high-quality database schemas. This module performs multi-path retrieval of tables, columns, and values based on the inputs and subsequently employs a model for iterative column selection.
@@
To realize this “society of experts” [20] philosophy, we propose a novel generation method that integrates our diverse ensemble of generators. At its core are multiple fine-tuned models, for which we capitalize on the high controllability of supervised fine-tuning (SFT) to construct multiple SQL generation models. Specifically, inspired by the multi-tasking approach in natural language processing, we explore a multi-task joint fine-tuning strategy tailored for SQL generation. This strategy involves the incorporation of various SQL and natural language transformation tasks to enhance the model’s syntactic alignment capabilities. Additionally, we investigate a multi-format SQL enhancement training approach designed to improve diversity. This involves fine-tuning specialized models to preferentially generate semantically equivalent queries that differ in structure or style, for instance, using complex Common Table Expressions (CTEs) versus simpler subqueries, or enforcing specific stylistic conventions. Ultimately, these strategies significantly improve the performance of fine-tuned models.
@@
For the SQL candidate selection phase, we employ a selection model that works in concert with a candidate reorganization strategy. The strategy first reorganizes candidates based on consistency results to enhance the model’s attention towards potentially correct options. Subsequently, the selection model, fine-tuned on constructed comparative samples, identifies the best answer from this structured list, leading to improved selection performance.
@@
In the experiments, we evaluate XiYan-SQL on two benchmarks, BIRD [13], and Spider [42], to demonstrate the superiority of the framework. For the well-known and challenging BIRD benchmark, we achieve an execution accuracy of 75.63%. This result established a new SOTA performance on the BIRD leaderboard upon our submission on December 17, 2024. For the Spider test set, we achieve SOTA performance with an execution accuracy of 89.65%. The extensive experimental results and comprehensive analysis further confirm the effectiveness and robustness of our method. The related code and models are being incrementally released through our project’s main repository to support community research.  [^2] https://github.com/XGenerationLab/XiYan-SQL
Our main contributions are as follows.
@@
  • We propose a novel selection approach, combining a candidate reorganization strategy with a powerful SQL selection model to identify the optimal SQL from multiple candidates.
@@ II Related Work
Prompt Engineering Methods.  Prompt engineering methods leverage the inherent capabilities of models to guide LLMs in generating diverse SQL queries by optimizing prompts [4, 5, 22, 23]. These methods are typically employed out-of-the-box on closed-source models with a vast number of parameters, such as GPT-4[1] and Gemini[29]. Some approaches based on in-context learning address various stages of the workflow, including schema linking[28, 19], SQL generation [9], and SQL refinement[23, 33, 38], by designing question/task decomposition, demonstration selection[8], chain-of-thought reasoning[36], and multi-agent processing within the prompts [33]. Regarding schema linking, a common objective in prior work is to prune the database schema to a single, most-relevant subset. This includes modern LLM-based methods (e.g., CHESS [28], CHASE-SQL [22], OpenSearch-SQL [37]) and earlier sophisticated systems that use dedicated classifiers (e.g., RESDSQL [10]). Even within the fine-tuning paradigm, methods like CodeS [11] tackle schema linking, albeit by integrating it into a monolithic model. While their techniques vary, their goal is a single optimized schema for a single generator. In stark contrast, our Schema Filter’s objective is to iteratively generate a diverse set of complementary schemas (e.g., a high-precision and a high-recall schema). This diversity is purpose-built to provide robust inputs for our downstream multiple SQL generation module.
To further enhance performance, an effective strategy is to generate multiple candidate SQL queries through different prompt designs and subsequently select the best one [8, 22]. However, these methods rely solely on prompt design, which presents challenges related to the sensitivity of LLMs to prompts, potentially compromising the quality and robustness of SQL generation, along with incurring significant inference overhead. In contrast, our approach effectively mitigates these problems by constructing different high-accuracy fine-tuned SQL generation models.
@@
While sharing high-level goals with the works discussed, XiYan-SQL distinguishes itself through its unique architectural philosophy and its relationship with the emerging agent paradigm. Architecturally, unlike prompt-based methods that query a single API model multiple times, or prior fine-tuning works that focus on improving a single model, we are the first to propose and systematically evaluate a hybrid framework centered on an ensemble of distinct, fine-tuned local models. This approach is not merely an assembly of parts, but a synergistic co-design: our iterative schema filter is built to produce diverse schemas for the multiple generators, and our novel multi-task/multi-format fine-tuning strategy is specifically created to induce the expert “personalities” required for the ensemble’s success. In the context of agent-based systems, our framework’s fully decoupled and modular design allows its components to be seamlessly embedded as powerful “tools” or “experts.” For instance, a general-purpose agent could delegate complex database interaction tasks to our entire pipeline. Thus, XiYan-SQL is positioned not as an alternative to the agent paradigm, but as a complementary and enabling technology for it. By demonstrating the power of this integrated, multi-generator architecture, our work presents a novel and practical blueprint for building robust, private, and cost-effective Text-to-SQL systems that can both stand alone and enhance future agent-based workflows.
@@ III-A Overall Framework
We present the XiYan-SQL framework in Figure 1. The inputs are: (i) the user question; (ii) external textual knowledge (hereafter “evidence”), provided in the form of short text snippets such as schema explanations, synonym mappings, or business rules; and (iii) the database schema (tables, columns, and values). The framework consists of three steps: Schema Filter, Multiple SQL Generation, and SQL Selection. The Schema Filter module initially generates multiple filtered schemas. Subsequently, multiple candidate SQL queries are generated through various SQL generators. Finally, the most accurate SQL query is selected by the selection model based on the candidate reorganization. The details of each step are described below.
@@ III-B Schema Filter
Multi-path Retrieval. Multi-path Retrieval is a pruning mechanism for large databases that employs cosine similarity for the embeddings of all tables, columns, and values. Initially, we used an out-of-the-box LLM $\mathcal{M}_{s}$ to extract keywords from the question $Q$ and evidence $E$, represented as $K=f_{\mathcal{M}_{s}}(Q,E)$. These keywords, denoted as $K=\{k_{1},k_{2},\ldots\}$, are instrumental in identifying relevant columns and values within the database. In the table and column retrieval stage, we compute the scores between the keywords $K$ and the original schema $S$, which can be represented as a set of columns, $\{c_{1},c_{2},\ldots\}$. This process is calculated as the product of two distinct scores as follows.
@@
Column Selection. Column Selection introduces diversity into the retrieved schemas, facilitating subsequent multiple SQL generation. This process is detailed in Algorithm 1, which outputs $p_{s}$ different schemas through multiple iterations. In each iteration, we prompt an LLM $\mathcal{M}_{s}$ to select columns related to the question and evidence, saved as $S^{slct}_{i}$. The function PFKeyIdentifier then identifies primary and foreign keys according to the columns in $S^{slct}_{i}$. It is designed to handle multi-column keys by leveraging the database’s metadata: it automatically completes partial keys (i.e., including all columns of a composite key if one is selected) and includes the corresponding referenced keys from other tables to preserve valid join paths. A new schema is generated by unifying the previous schemas with the selected columns. We update $S^{rtrv}$ by removing only the selected non-key columns and explicitly retaining key columns (primary/foreign keys) to maintain schema integrity and connectivity. As iterations increase, the precision of the selected schemas tends to decrease while recall improves, resulting in a diversity of filtered schemas.
@@
7:  $S^{rtrv}\leftarrow S^{rtrv}\setminus(S^{slct}_{i}\setminus P_{i})$ {Remove selected non-key columns, preserve PK/FK.}
@@ III-C Multiple SQL Generation
However, this unidirectional training (Text-to-SQL) is insufficient to build a deep, bidirectional understanding between language and SQL. To address this issue and enhance the semantic alignment capability of LLMs for this task, we further design auxiliary alignment tasks and explicitly fine-tune the SQL generation model for execution. In response to natural language questions, we design a reverse inference task, as illustrated in Figure 2(b), which aims to encourage the model to infer a set of potential questions based on SQL and relevant information. For natural language evidence knowledge, we establish a task focused on the reverse inference of evidence, shown in Figure 2(c). This task involves the model identifying the most relevant evidence from a set of candidate evidence based on the SQL generation process. Additionally, for SQL queries, we develop a self-refine task for the model, as depicted in Figure 2(d), which entails regenerating SQL after refinement based on previous SQL generation processes and execution results. By leveraging a series of specialized tasks to jointly fine-tune models, our developed single generator surpasses ICL-based closed-source models, such as GPT-4o and Gemini, resulting in the generation of higher-quality SQL queries. The detailed data construction methods for these auxiliary tasks are provided in Appendix II.B to ensure reproducibility.
Figure 2: Illustration of the four tasks in our Multi-Task Fine-Tuning strategy. It shows the input-output transformations for (a) the standard Text-to-SQL task, and the auxiliary tasks of (b) Question Inference, (c) Evidence Inference, and (d) Self-Refine.
We further develop from a single generator to multiple generators to enhance the diversity of generated candidates. For a single user question, there can be multiple SQL queries that correspond to fulfill the intent. Therefore, we aim to develop additional models with different ways of generation to increase the likelihood of arriving at the correct answers through diverse outputs. To achieve this, we introduce “multi-format” SQL variations during model construction. These added reformulations fall into two primary types: (1) Structural Variations: Encouraging the model to use more complex structures (e.g., rewriting a subquery as a Common Table Expression (CTE)), as illustrated by $\mathrm{SQL}_{1}$ in Figure 3. (2) Stylistic Variations: Adopting different writing conventions (e.g., standardizing keyword casing and table aliases), as shown by $\mathrm{SQL}_{3}$ in Figure 3. These multi-format data are then used to train specialized generators, with the construction methods for these diverse variations detailed in Appendix II.B. By leveraging these formatted and multi-task data for joint fine-tuning, we obtain high-accuracy models with diverse SQL generation preferences.
@@
Figure 3: Examples of “multi-format” SQL queries corresponding to the same input. Compared to the typical response (middle), the left query demonstrates a Structural Variation (e.g., using a more complex structure), while the right query illustrates a Stylistic Variation (e.g., adopting a different writing convention).
@@ III-D SQL Selection
After obtaining multiple SQL candidates, the challenge remains to select the most accurate and reasonable SQL query from these candidates. Most candidate selection methods [8, 28] employ self-consistency [34] to select the SQL query that appears most consistently across multiple candidate samples. However, there are limitations present: it cannot adequately handle situations where there is a lack of majority consensus, and even the most consistent result is not always the correct case. To this end, we propose a model-based selection approach, which consists of a selection model and a candidate reorganization strategy to guide its attention.
To develop an effective selection model, we focus on constructing a rich listwise comparison dataset. This process begins with generating diverse candidate SQL queries using multiple generators, which include our ensemble of fine-tuned models and powerful ICL-based models (e.g., GPT-4o). With these candidates, we construct training instances composed of positive and negative samples. Recognizing that multiple queries can be semantically correct, we identify positive samples not just by matching the ground-truth SQL, but by verifying any candidate that produces an identical execution result. Negative samples are obtained through a multi-pronged strategy, involving not only natural failures (e.g., syntax, execution, or logical errors) but also more challenging “near-miss” examples created via LLM-based “controlled modifications”—a process where we introduce subtle, deliberate errors into correct queries. Before final dataset assembly, all candidates undergo a “de-formalization process”: we apply a standard SQL formatter to remove superficial stylistic differences (e.g., casing and spacing), forcing the model to learn semantic correctness rather than spurious generator-specific patterns. A comprehensive breakdown of this entire data construction pipeline is provided in Appendix II.C to ensure reproducibility.
During inference, to help the model better handle the numerous candidates, we apply the aforementioned candidate reorganization strategy to enhance its attention. To align with the training phase, all candidates presented to the selection model are first de-formalized to remove superficial stylistic artifacts. The complete selection process is illustrated in Algorithm 3. Initially, we obtain the execution results $R=\{r_{1},\ldots,r_{n}\}$ of the candidate SQL queries $L=\{l_{1},\ldots,l_{n}\}$ on the database, ensuring that any candidates with execution errors have already been excluded. We group the candidates based on the consistency of their execution results to obtain a collection of clusters $\mathcal{C}=\{C_{1},...,C_{m}\}$. Subsequently, we perform inter-group and intra-group sorting. The inter-group sorting is performed in descending order based on the size of each group, resulting in $\mathcal{C}^{\prime}$. The intra-group sorting is executed according to the order of the generators $\mathcal{O}$ corresponding to the candidates, resulting in $\mathcal{C}^{\prime\prime}$, where the sequence $\mathcal{O}$ is arranged in descending order based on the performance evaluation of the generators.
@@
[…]
2: Clustering $\mathcal{C}=\{C_{1},...,C_{m}\}\leftarrow\text{groupby}(L,R)$
3: Schema $\mathcal{S}^{un}\leftarrow\text{schema\_union}(\mathcal{S})$
4:if $|\mathcal{C}|=1$ then
5:  return $l^{*}\leftarrow\arg_{l\in L}\min|l|$ {All are consistent; select shortest as a heuristic for simplicity.}
6:else
7:  Clusters $\mathcal{C}^{\prime}\leftarrow\text{sort}(\mathcal{C})$ by size, descending {Inter-group sorting}
8:  Initialize Clusters $\mathcal{C}^{\prime\prime}\leftarrow\emptyset$
9:  for each cluster $C^{\prime}_{i}\in\mathcal{C}^{\prime}$ do
10:   Append $C^{\prime\prime}_{i}\leftarrow\text{sort}(C^{\prime}_{i})$ order by $\mathcal{O}$ {Intra-group sorting}
11:  Initialize Candidate $L^{\prime}\leftarrow\emptyset$, indicator $j\leftarrow 0$
12:  if $|C_{1}^{\prime\prime}|\geq\lceil{|L|}/2\rceil$ then
13:   for each Cluster $C_{i}^{\prime\prime}\in\mathcal{C}^{\prime\prime}$ do
14:    for each SQL $l\in C_{i}^{\prime\prime}$ do
15:     Append $L_{j}^{\prime}\leftarrow l$ , $j\leftarrow j+1$
16:  else
17:   for each Cluster $C_{i}^{\prime\prime}\in\mathcal{C}^{\prime\prime}$ do
18:    Append $L_{j}^{\prime}\leftarrow\arg_{l\in C_{i}^{\prime\prime}}\min|l|$, $j\leftarrow j+1$
19:  return $l^{*}\leftarrow f_{\mathcal{M}_{c}}(Q,\mathcal{S}^{un},E,L^{\prime})$ {Selection model prediction}
If there exists a consensus result that predominates within the clusters, namely $|C_{1}^{\prime\prime}|\geq\lceil{|L|}/2\rceil$, we organize all candidates sequentially according to order $\mathcal{C}^{\prime\prime}$ and present them to the selection model $\mathcal{M}_{c}$. Conversely, we select one candidate from each group within $\mathcal{C}^{\prime\prime}$ and assemble the results in accordance with the inter-group ordering to present to $\mathcal{M}_{c}$. The model $\mathcal{M}_{c}$ takes into account the question, the union of all filtered schemas, the evidence, and the reorganized candidates to make its decision. Due to the selection tendency of LLMs towards options that appear earlier [45], employing this clustering and ranking approach to organize candidates allows us to effectively leverage strong consistency priors while mitigating the influence of interfering information. This allows the model’s attention to be concentrated, thereby enhancing its comprehension ability.